# DRUPI: Dataset Reduction Using Privileged Information

### v2

@@ Abstract
Dataset reduction (DR) seeks to select or distill samples from large datasets into smaller subsets while preserving performance on target tasks. Existing methods primarily focus on pruning or synthesizing data in the same format as the original dataset, typically the input data and corresponding labels. However, in DR settings, we find it is possible to synthesize more information beyond the data-label pair as an additional learning target to facilitate model training. In this paper, we introduce Dataset Reduction Using Privileged Information (DRUPI), which enriches DR by synthesizing privileged information alongside the reduced dataset. This privileged information can take the form of feature labels or attention labels, providing auxiliary supervision to improve model learning. Our findings reveal that effective feature labels must balance between being overly discriminative and excessively diverse, with a moderate level proving optimal for improving the reduced dataset’s efficacy. Extensive experiments on ImageNet, CIFAR-10/100, and Tiny ImageNet demonstrate that DRUPI integrates seamlessly with existing dataset reduction methods, offering significant performance gains. The code will be released after the paper is accepted.
@@ 1 Introduction
Dataset Reduction (DR) has attracted considerable attention in recent years, with the primary aim of compressing large datasets into smaller subsets while maintaining comparable statistical performance. Existing methods for DR can be broadly classified into two main categories: coreset selection and dataset distillation. Coreset selection methods focus on selecting a subset of samples from the original dataset (Har-Peled & Mazumdar, 2004; Welling, 2009; Toneva et al., 2018), while dataset distillation involves synthesizing unseen samples from the dataset (Wang et al., 2018; Zhao et al., 2020; Zhao & Bilen, 2022; Cazenavette et al., 2022; Wang et al., 2024).
@@
In typical real-world scenarios, training models for target tasks is generally constrained to input data (e.g., images) and their corresponding labels, as these are the most readily available elements. While existing DR methods have shown strong performance (Wang et al., 2018; Zhao et al., 2020; Zhao & Bilen, 2022; Cazenavette et al., 2022; Yin et al., 2023), they typically do so by compressing datasets in the same or similar format, such as the conventional data-label structure. Even advanced dataset distillation techniques, which re-parameterize images or labels to create alternative representations (Kim et al., 2022; Zhao et al., 2023; Deng & Russakovsky, 2022; Liu et al., 2022; Wei et al., 2024), are limited by this conventional framework. As illustrated in Figure 1(a), this reliance on fixed data-label structures restricts the capacity of such methods to incorporate richer information that could further enhance model training and improve generalization.
@@
In fact, DR settings offer the potential to create more diverse compressed datasets that extend beyond simple input data $x_{i}$ and labels $y_{i}$, incorporating richer forms of information. A notable example is the concept of privileged information, first introduced in the context of statistical learning (Vapnik & Vashist, 2009; Pechyony & Vapnik, 2010). Figure 2 provides a illustration for privileged information. Let us consider a more concrete example, where $x_{i}$ might represent a biopsy image, and the privileged information $f_{i}^{\star}$ for $x_{i}$ could be the oncologist’s written assessment of the image (Vapnik et al., 2015). The label $y_{i}$ would then indicate whether the tissue in the image is malignant or benign. By leveraging this privileged information $f_{i}^{\star}$, a medical expert can make more informed decisions, benefiting from additional insights that improve diagnostic accuracy.
@@
However, none of the existing methods compress the original dataset beyond the traditional data-label structure or synthesize privileged information for auxiliary supervision. To address this gap, we introduce a novel approach that, for the first time, incorporates not only images and labels but also privileged information. Our method, called Dataset Reduction Using Privileged Information (DRUPI), is illustrated in Figure 1(b). We primarily synthesize feature labels for the reduced dataset, as these labels capture richer, high-dimensional information, enhancing dataset quality. These feature labels generalize effectively across various neural network architectures and provide a unified representation of latent statistics across multiple models, offering additional supervision during training. Additionally, we propose a more efficient form of feature labels, i.e., attention labels. As shown in Figure 1(c), the incorporation of privileged information produces gradients more aligned with those of the original dataset, ultimately improving the model’s generalization capabilities.
@@
Figure 1: A comparison between conventional dataset reduction pipelines and our proposed DRUPI framework. (a) Previous dataset reduction methods distill or select a subset $\mathcal{D}_{\mathcal{S}}$ from the original dataset $\mathcal{D}_{\mathcal{T}}$, maintaining the original “data-label” structure. (b) In contrast, DRUPI synthesizes auxiliary privileged information from $\mathcal{D}_{\mathcal{T}}$, enriching further supervision to models trained on the reduced subset $\mathcal{D}_{\mathcal{S}}$. (c) Cosine similarity between the gradients of a pre-trained model on synthetic datasets w/ and w/o privileged information (feature labels) and the real dataset. Synthetic datasets are generated using DC with 10 IPC. We used the same pre-trained ConvNet for gradient extraction.
@@
A key challenge for DRUPI lies in determining appropriate feature labels to synthesize for the reduced dataset. To address this, we employ an arbitrary methods of dataset distillation to synthesize the feature labels. During each step of the bi-level optimization, we match the statistical information of models trained on reduced datasets with and without feature labels. Furthermore, we observed that synthesized feature labels cannot be overly discriminative or diverse, which degrade the overall quality of the reduced dataset. This finding suggests that an optimal balance between discriminability and diversity is crucial for synthesizing effective feature labels. Our contributions are summarized as follows:
@@
    We propose a new paradigm, i.e., DRUPI, for dataset reduction. In particular, privileged information, such as feature labels, can be synthesized in addition to traditional data-label pairs. This privileged information provides additional generative supervision during model training, thereby improving the generalization ability of the reduced dataset.
@@
    We further provide a theoretical analysis of our DRUPI pipeline based on VC theory (Vapnik, 1998) from statistical learning, which rigorously ensures its effectiveness.
@@
    Our experiments demonstrate that DRUPI can be seamlessly integrated into state-of-the-art DR methods. Particularly, for coreset selection methods, applying DRUPI to Herding on CIFAR10 (with a fraction of 0.4%) improves performance by 24.3%, while applying it to K-center in cross-architecture evaluations leads to an improvement of up to 23.4%. For dataset distillation methods, integrating DRUPI with the DC method on CIFAR100 (with 10 images per class) yields a 4% improvement, and further cross-architecture evaluations of DC show gains of up to 18.3%.
2 Background and Related Work
@@
Similar to prior works on dataset reduction (DR), we focus on classification tasks (Welling, 2009; Zhao et al., 2020; Shin et al., 2023). Consider a multi-classification problem. Let $\mathcal{X}\in\mathbb{R}^{d}$ represent the input space and $\mathcal{Y}$ denote the set of possible labels. Our dataset, $\mathcal{D}_{\mathcal{T}}=\{(x_{i},y_{i})\}_{i=1}^{n}\subseteq\mathcal{X}\times\mathcal{Y}$, consists of $n$ training samples, where each $x_{i}\in\mathcal{X}$ is an input vector and $y_{i}\in\mathcal{Y}$ is its corresponding label. In typical DR settings, the goal is to obtain a smaller dataset, $\mathcal{D}_{\mathcal{S}}$ with size $m$, where $m\ll n$. DR methods are typically divided into two categories: coreset selection, where the reduced dataset is a subset of the original, and dataset distillation, where the reduced dataset consists of synthesized data not present in the original set but learned through optimization.
@@
Coreset selection techniques. Coreset selection techniques are designed to identify a representative subset $\mathcal{D}_{\mathcal{S}}$ from the complete dataset $\mathcal{D}_{\mathcal{T}}$. This process typically seeks to optimize a criterion that quantifies the informativeness of $\mathcal{D}_{\mathcal{S}}$, which matches that of the original dataset $\mathcal{D}_{\mathcal{T}}$. The informativeness can be gauged by various metrics, including gradients (Paul et al., 2021; Mirzasoleiman et al., 2020; Killamsetty et al., 2021a), loss values (Toneva et al., 2018), predictive uncertainties (Coleman et al., 2019b), proximity to decision boundaries (Ducoffe & Precioso, 2018; Margatina et al., 2021), and the sharpness of the learned model (Shin et al., 2023).
Dataset distillation methods. Dataset distillation approaches offer an alternative strategy by focusing on synthesizing a distilled dataset $\mathcal{D}_{\mathcal{S}}$ from the original dataset $\mathcal{D}_{\mathcal{T}}$, rather than directly selecting a subset. This is typically accomplished through a bi-level optimization process that aligns the performance of $\mathcal{D}_{\mathcal{S}}$ with that of $\mathcal{D}_{\mathcal{T}}$. A distance metric $\mathbf{D}$ is employed to quantify the statistical divergence between datasets, guiding the learning of $\mathcal{D}_{\mathcal{S}}$ via gradient descent. Specifically, the distilled dataset is updated as follows: $\mathcal{D}_{\mathcal{S}}\leftarrow\mathcal{D}_{\mathcal{S}}-\eta\cdot\nabla_{\mathcal{D}_{\mathcal{S}}}\mathbf{D}(\mathcal{D}_{\mathcal{S}},\mathcal{D}_{\mathcal{T}})$. The choice of distance metric $\mathbf{D}$ is versatile and can encompass various aspects such as gradients (Zhao et al., 2020; Lee et al., 2022; Zhao & Bilen, 2021), feature representations (Zhao & Bilen, 2022; Sajedi et al., 2023), training trajectories (Cazenavette et al., 2022; Du et al., 2023; Cui et al., 2023; Guo et al., 2023), and kernel information (Nguyen et al., 2020; 2021; Zhou et al., 2022).
@@
In addition to direct performance matching, certain methodologies endeavor to re-parameterize input data to enhance compression efficiency. Techniques employed in this context include exploiting data regularity, as discussed in various studies (Kim et al., 2022; Zhao et al., 2023; Son et al., 2024), factorizing images to capture intrinsic structures (Liu et al., 2022; Deng & Russakovsky, 2022), and employing sparse coding to represent data effectively (Wei et al., 2024).
3 Method
@@ 3.1 Determining Privileged Information
Although prior research on dataset reduction has shown impressive results, it mainly focuses on generating reduced datasets in conventional data-label formats, as depicted in Figure 2(a). However, more informative data, such as privileged information, can be utilized to enhance both the utility and representational performance of reduced datasets, as illustrated in Figure 2(b). Below, we briefly explore several forms of privileged information that can be incorporated to achieve this goal.
Soft labels. We first claim that soft labels are a form of privileged information, as they offer richer insight into how an expert model interprets predictions, providing soft probabilities rather than a single hard label. Specifically, the non-target logits can be viewed as additional information for supervision. Several works have previously explored the effectiveness of soft labels in dataset reduction (Guo et al., 2023; Cui et al., 2023; Bohdal et al., 2020; Qin et al., 2024). However, while soft labels enhance the available information, they are limited to low-dimensional discriminative representations and fail to capture more complex, high-dimensional statistics. Moreover, they do not fundamentally alter the data-label structure of the reduced dataset representation.
@@
Attention labels. Alongside feature labels, we propose attention labels as an alternative form of privileged information that provides a more memory-efficient representation. Attention labels can be derived from either spatial or channel attention of feature labels (Woo et al., 2018). For example, given a feature label of size $Ch\times H\times W$, spatial attention reduces the $Ch$ dimensions through pooling operations (e.g., average pooling, max pooling), resulting in an attention label of size $1\times H\times W$. Similarly, channel attention applies pooling along the $H\times W$ dimension to produce a reduced representation, i.e., $Ch\times 1\times 1$. Both feature and attention labels are valuable, with attention labels offering a more efficient representation but with a possible trade-off in the richness of information. We provide more discussion on attention labels in Appendix B.2.
In this work, we primarily focus on generating additional feature labels [^1] We consider attention labels as a specific form of feature labels. Therefore, for simplicity, we use the term “feature labels” as a unified description for both feature labels and attention labels. for the reduced dataset, as these forms of privileged information provide more complementary and useful insights for model training. However, privileged information can take various forms beyond attention and feature labels. Depending on the task and the model architecture, other types of information, such as learned embeddings, domain-specific signals, or task-related metadata, could be equally beneficial in enhancing the informativeness and performance of reduced datasets. The flexibility to incorporate different kinds of privileged information allows us to tailor the dataset to specific needs and maximize the potential of the reduced data.
@@ 3.2 Synthesizing Privileged Information
Learning Feature Labels. A more robust approach to obtaining feature labels is through learning-based methods. Many dataset distillation techniques can be adapted for learning feature labels. For instance, we can employ the Dataset Condensation (DC) (Zhao et al., 2020) method as an illustrative example to guide the process of learning synthetic feature labels. Suppose we have a learned synthetic dataset $\mathcal{D}_{\mathcal{S}}$. In the typical DC method, we initialize a random model $g$ with parameters $\theta=\theta_{0}$ and train it for $T$ epochs on both $\mathcal{D}_{\mathcal{T}}$ and $\mathcal{D}_{\mathcal{S}}$ separately. The synthetic dataset $\mathcal{D}_{\mathcal{S}}$ is updated by matching the category gradients between the two datasets, which can be expressed as follows [^2] We ignore the category symbol for simplicity.:
@@
$\mathcal{L}_{cls}=\underset{(\tilde{x}_{i},\tilde{y}_{i})\in\mathcal{D}_{\mathcal{S}}^{\star}}{\mathbb{E}}[\ell_{ce}\left(y_{i},\sigma(g(\tilde{x}_{i};\theta_{t}))\right)],$
@@
where $\lambda_{reg}$ is a hyper-parameter to determine the scale of using privileged information. In addition to DC, other dataset distillation methods can also be employed to synthesize feature labels $f_{i}^{\star}$. We provide further resuls on coreset selection methods like Herding (Welling, 2009), K-center (Har-Peled & Mazumdar, 2004), Forgetting (Toneva et al., 2018), and dataset distillation methods like DC (Zhao et al., 2020), MTT (Cazenavette et al., 2022), and DATM (Guo et al., 2023).
In addtional, we introduce additional supervision to enhance the discriminative power of these feature labels while preserving their diversity.
@@
$\bullet$ Versatility synthesization. Beyond task-specific supervision, we aim to preserve the generative versatility of the feature labels, i.e., a set of feature labels $f_{i}^{\star}\in F_{i}$, where $F_{i}$ represents the possible feature label set for input $x_{i}$. This approach involves synthesizing multiple feature labels for a single data-label pair. Appropriate versatility enhancement ensures that the synthesized feature labels remain informative across different tasks and applications, providing a richer and more comprehensive representation of the data. When using multiple feature labels, we primarily employ two strategies: randomly selecting one or using the average feature label from $F_{i}$. Further discussions demonstrating the benefits of this versatility are presented in Figure 4(a) and Appendix B.3.
@@
where $F_{i}$ denotes the feature label set, containing multiple $f_{i}^{\star}$ for a single data-label pair $(\tilde{x}_{i},\tilde{y}_{i})$. The pseudocode for learning privileged information is provided in Appendix D.

### v3

# DCPI: Dataset Condensation using Privileged Information

@@ Abstract
Dataset Condensation (DC) seeks to select or distill samples from large datasets into smaller subsets while preserving performance on target tasks. Existing methods primarily focus on pruning or synthesizing data in the same format as the original dataset, typically being the input data and corresponding labels. However, in DC settings, we find it is possible to synthesize more information beyond the data-label pair as an additional learning target to facilitate model training. In this paper, we introduce Dataset Condensation using Privileged Information (DCPI), which enriches DC by synthesizing privileged information alongside the reduced dataset. This privileged information can take the form of feature labels or attention labels, providing auxiliary supervision to improve model learning. Our findings reveal that effective feature labels must balance between being overly discriminative and excessively diverse, with a moderate level proves optimal for improving the reduced dataset’s efficacy. Extensive experiments on ImageNet-1K, CIFAR-10/100 and Tiny ImageNet demonstrate that DCPI integrates seamlessly with existing dataset condensation methods, offering significant performance gains.
@@ 1 Introduction
Dataset Condensation (DC) has attracted considerable attention in recent years, with the primary aim of compressing large datasets into smaller subsets while maintaining comparable statistical performance. The existing DC methods can be broadly classified into two main categories: coreset selection and dataset distillation. Coreset selection methods focus on selecting a subset of samples from the original dataset Har-Peled and Mazumdar (2004); Welling (2009); Toneva et al. (2018), while dataset distillation involves synthesizing unseen samples from the dataset Wang et al. (2018); Zhao et al. (2020); Zhao and Bilen (2022); Cazenavette et al. (2022); Wang et al. (2024).
@@
In typical real-world scenarios, training of models for target tasks are generally restricted to input data (e.g., images) and their corresponding labels, as these are the most readily available elements. Although existing DC methods have shown strong performance Wang et al. (2018); Zhao et al. (2020); Zhao and Bilen (2022); Cazenavette et al. (2022); Yin et al. (2023), they typically do so by compressing datasets in the same or similar format, such as the conventional data-label structure. Even advanced dataset distillation techniques, which re-parameterize images or labels to create alternative representations Kim et al. (2022); Zhao et al. (2023); Deng and Russakovsky (2022); Liu et al. (2022); Wei et al. (2024), are restricted by this conventional framework. As illustrated in Figure 1(a), this reliance on fixed data-label structures restricts the ability of such methods to incorporate richer information that could further enhance model training and improve generalization.
@@
In fact, DC settings offer the potential to create more diverse compressed datasets that extend beyond the simple input data $x_{i}$ and labels $y_{i}$, incorporating richer forms of information. A notable example is the concept of privileged information, first introduced in the context of statistical learning Vapnik and Vashist (2009); Pechyony and Vapnik (2010); Wang et al. (2026). Figure 2 provides an illustration of the privileged information. Let us consider a more concrete example, where $x_{i}$ might represent a biopsy image, and the privileged information $f_{i}^{\star}$ for $x_{i}$ could be the oncologist’s written assessment of the image Vapnik et al. (2015). The label $y_{i}$ would then indicate whether the tissue in the image is malignant or benign. By leveraging this privileged information $f_{i}^{\star}$, a medical expert can make more informed decisions, benefiting from additional insights that improve diagnostic accuracy.
@@
However, none of the existing methods compress the original dataset beyond the traditional data-label structure or synthesize privileged information for auxiliary supervision. To address this gap, we introduce a novel approach that, for the first time, incorporates not only images and labels but also privileged information. Our method, called Dataset Condensation using Privileged Information (DCPI), is illustrated in Figure 1(b). We primarily synthesize feature labels for the reduced dataset, as these labels capture richer, high-dimensional information, enhancing dataset quality. These feature labels generalize effectively across various neural network architectures and provide a unified representation of latent statistics across multiple models, offering additional supervision during training. Additionally, we propose a more efficient form of feature labels, i.e., attention labels. As shown in Figure 1(c), the incorporation of privileged information produces gradients more aligned with those of the original dataset, ultimately improving the generalization capabilities of the model.
@@
Figure 1: A comparison between conventional dataset reduction pipelines and our proposed DCPI framework. (a) Previous dataset reduction methods distill or select a subset $\mathcal{D}_{\mathcal{S}}$ from the original dataset $\mathcal{D}_{\mathcal{T}}$, maintaining the original “data-label” structure. (b) In contrast, DCPI synthesizes auxiliary privileged information from $\mathcal{D}_{\mathcal{T}}$, enriching further supervision to models trained on the reduced subset $\mathcal{D}_{\mathcal{S}}$. (c) Cosine similarity between the gradients of a pre-trained model on synthetic datasets w/ and w/o privileged information (feature labels) and the real dataset. Synthetic datasets are generated using DC with 10 images per class (IPC) on the CIFAR-10 and CIFAR-100 datasets. We used the same pre-trained ConvNet for gradient extraction.
@@
A key challenge for DCPI lies in determining the appropriate feature labels to synthesize for the reduced dataset. To address this, we employ an arbitrary method of dataset distillation to synthesize the feature labels. During each step of the bi-level optimization, we match the statistical information of models trained on reduced datasets with and without feature labels. Furthermore, we observed that synthesized feature labels cannot be overly discriminative or diverse, which degrades the overall quality of the reduced dataset. This finding suggests that an optimal balance between discriminability and diversity is crucial for synthesizing effective feature labels. Our contributions are summarized as follows:
@@
    We propose a new paradigm, i.e., DCPI, for dataset reduction. In particular, privileged information, such as feature labels, can be synthesized in addition to traditional data-label pairs. This privileged information provides additional generative supervision during model training, thereby improving the generalization ability of the reduced dataset.
@@
    We further provide a theoretical analysis of our DCPI pipeline based on VC theory Vapnik (1998) from the statistical learning, which rigorously ensures its effectiveness.
@@
    Our experiments demonstrate that DCPI can be seamlessly integrated into state-of-the-art DC methods. Particularly, for coreset selection methods, applying DCPI to Herding on CIFAR10 (with a fraction of 0.4%) improves performance by 24.3%, while applying it to K-center in cross-architecture evaluations leads to an improvement of up to 23.4%. For dataset distillation methods, integrating DCPI with the DC method on CIFAR100 (with a fraction of 0.2%) yields a 2.1% improvement, and further cross-architecture evaluations of DC show gains of up to 18.3%.
2 Related Work
@@
Following prior work on dataset condensation (DC) Welling (2009); Zhao et al. (2020); Shin et al. (2023), we consider a multi-classification problem. Let $\mathcal{X}\in\mathbb{R}^{d}$ be the input space and $\mathcal{Y}$ the label set. The dataset $\mathcal{D}_{\mathcal{T}}=\{(x_{i},y_{i})\}_{i=1}^{n}\subseteq\mathcal{X}\times\mathcal{Y}$ has $n$ samples, with $x_{i}\in\mathcal{X}$ as input vectors and $y_{i}\in\mathcal{Y}$ as labels. DR aims to create a smaller dataset $\mathcal{D}_{\mathcal{S}}$ of size $m\ll n$. DR methods include coreset selection, using a subset of the original data, and dataset distillation, using optimized synthetic data not in the original set.
@@
Coreset Selection Techniques. Coreset selection identifies a representative subset $\mathcal{D}_{\mathcal{S}}$ from dataset $\mathcal{D}_{\mathcal{T}}$, optimizing a criterion for informativeness matching $\mathcal{D}_{\mathcal{T}}$. Informativeness is measured by metrics like gradients Paul et al. (2021); Mirzasoleiman et al. (2020); Killamsetty et al. (2021a), loss values Toneva et al. (2018), predictive uncertainties Coleman et al. (2019), decision boundary proximity Ducoffe and Precioso (2018); Margatina et al. (2021), or model sharpness Shin et al. (2023).
Dataset Distillation Methods. Dataset distillation synthesizes a distilled dataset $\mathcal{D}_{\mathcal{S}}$ from $\mathcal{D}_{\mathcal{T}}$ via bi-level optimization to match $\mathcal{D}_{\mathcal{T}}$’s performance. A distance metric $\mathbf{D}$ measures statistical divergence, guiding updates: $\mathcal{D}_{\mathcal{S}}\leftarrow\mathcal{D}_{\mathcal{S}}-\eta\cdot\nabla_{\mathcal{D}_{\mathcal{S}}}\mathbf{D}(\mathcal{D}_{\mathcal{S}},\mathcal{D}_{\mathcal{T}})$. Metric $\mathbf{D}$ may include gradients Zhao et al. (2020); Lee et al. (2022a); Zhao and Bilen (2021), features Zhao and Bilen (2022); Sajedi et al. (2023), training trajectories Cazenavette et al. (2022); Du et al. (2023); Cui et al. (2023); Guo et al. (2023), or kernel information Nguyen et al. (2020; 2021); Zhou et al. (2022).
@@
In addition to direct performance matching, certain methodologies endeavor to re-parameterize input data to enhance compression efficiency. Techniques employed in this context include exploiting data regularity, as discussed in various studies Kim et al. (2022); Zhao et al. (2023); Son et al. (2024), factorizing images to capture intrinsic structures Liu et al. (2022); Deng and Russakovsky (2022), and employing sparse coding to represent data effectively Wei et al. (2024).
3 Methodology
@@ 3.1 Determining Privileged Information
Although prior research on dataset condensation has shown impressive results, it mainly focuses on generating reduced datasets in conventional data-label formats, as depicted in Figure 2(a). However, more informative data, such as privileged information, can be utilized to enhance both the utility and representational performance of reduced datasets, as illustrated in Figure 2(b). Below, we briefly explore several forms of privileged information that can be incorporated to achieve this goal.
Soft labels. We first claim that soft labels are a form of privileged information, as they offer richer insight into how an expert model interprets predictions, providing soft probabilities rather than a single hard label. Specifically, the non-target logits can be viewed as additional information for supervision. Several works have previously explored the effectiveness of soft labels in dataset condensation Guo et al. (2023); Cui et al. (2023); Bohdal et al. (2020); Qin et al. (2024). However, while soft labels enhance the available information, they are limited to low-dimensional discriminative representations and fail to capture more complex, high-dimensional statistics. Moreover, they do not alter the data-label structure of the reduced dataset representation.
@@
Attention labels. Alongside feature labels, we propose attention labels as an alternative form of privileged information that provides a more memory-efficient representation. Attention labels can be derived from either spatial or channel attention of feature labels Woo et al. (2018). For example, given a feature label of size $Ch\times H\times W$, spatial attention reduces the $Ch$ dimensions through pooling operations (e.g., average pooling, max pooling), resulting in an attention label of size $1\times H\times W$. Similarly, channel attention applies pooling along the $H\times W$ dimension to produce a reduced representation, i.e., $Ch\times 1\times 1$. Both feature and attention labels are valuable, with attention labels offering a more efficient representation but with a possible trade-off in the richness of information.
In this work, we primarily focus on generating additional feature labels [^1] We consider attention labels as a specific form of feature labels. Therefore, for simplicity, we use the term “feature labels” as a unified description for both feature labels and attention labels. for the reduced dataset, as these forms of privileged information provide more complementary and useful insights for model training. However, privileged information can take various forms beyond attention and feature labels. Depending on the task and the model architecture, other types of information, such as learned embeddings, domain-specific signals, or task-related metadata, could be equally beneficial in enhancing the informativeness and performance of reduced datasets. Incorporating diverse privileged information lets us tailor datasets to specific needs and maximize reduced data potential.
@@ 3.2 Synthesizing Privileged Information
Learning Feature Labels. A more robust approach to obtaining feature labels is through learning-based methods. Many dataset distillation techniques can be adapted for learning feature labels. For instance, we can employ the Dataset Condensation (DC) Zhao et al. (2020) method as an illustrative example to guide the process of learning synthetic feature labels. Suppose we have a learned synthetic dataset $\mathcal{D}_{\mathcal{S}}$. In the typical DC method, we initialize a random model $g$ with parameters $\theta=\theta_{0}$ and train it for $T$ epochs on both $\mathcal{D}_{\mathcal{T}}$ and $\mathcal{D}_{\mathcal{S}}$ separately. The synthetic dataset $\mathcal{D}_{\mathcal{S}}$ is updated by matching the category gradients between the two datasets, which can be expressed as follows [^2] We ignore the category symbol for simplicity.:
@@
$\mathcal{L}_{cls}=\underset{(\tilde{x}_{i},\tilde{y}_{i})\in\mathcal{D}_{\mathcal{S}}^{\star}}{\mathbb{E}}[\ell_{ce}\left(\tilde{y}_{i},\sigma(g(\tilde{x}_{i};\theta_{t}))\right)],$
@@
where $\lambda_{reg}$ is a hyper-parameter to determine the scale of using privileged information. In addition to DC, other dataset distillation methods can also be used synthesize feature labels $f_{i}^{\star}$. We provide further results on coreset selection methods like Herding Welling (2009), K-center Har-Peled and Mazumdar (2004), Forgetting Toneva et al. (2018), and dataset distillation methods like DC Zhao et al. (2020), MTT Cazenavette et al. (2022), and DATM Guo et al. (2023).
In addition, we introduce additional supervision to enhance the discriminative power of these feature labels while preserving their diversity.
@@
$\bullet$ Versatility synthesization. Beyond task-specific supervision, we aim to preserve the generative versatility of the feature labels, i.e., a set of feature labels $f_{i}^{\star}\in F_{i}$, where $F_{i}$ represents the possible feature label set for input $x_{i}$. This approach involves synthesizing multiple feature labels for a single data-label pair. Appropriate versatility enhancement ensures that the synthesized feature labels remain informative across different tasks and applications, providing a richer and more comprehensive representation of the data. When using multiple feature labels, we primarily employ two strategies: randomly selecting one or using the average feature label from $F_{i}$. Further discussions demonstrating the benefits of this versatility are presented in Figure 4(a) and Table 11 in Appendix B.2.
@@
where $F_{i}$ denotes the feature label set, containing multiple $f_{i}^{\star}$ for a single data-label pair $(\tilde{x}_{i},\tilde{y}_{i})$. The pseudocode for learning privileged information is provided in Algorithm 1 in Appendix D.